Operations Report – February 17th

SOL: 7

Name of person filing report: Verónica Triviño

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on at 1600 (17th Feb), turned off at 0830 (17th Feb)

Solar— SOC % – Turned off Gen (17th Feb) 82%, Turned on Gen (17th Feb) 100%

Diesel Reading – 90%

Propane Reading – ~ 40%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 62%, ~ 343 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 216.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01413047 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 51.0 hours

Beginning charge: 100%

Ending charge: 28%

Currently charging: Yes

Curiosity rover used: Used

Hours: 75.0 hours

Beginning charge: 100%

Ending charge: 44%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 75.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: N/A

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? N/A

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Some suits are being repaired in the evening. The parts of the ones repaired yesterday evening worked well on this morning’s EVA.

Summary of Hab operations: Screws tightened in the back airlock door. White tunnel covers were re-secured with new zip ties, as well as tightening loose zipties and clipping the ends of the zip ties.

Summary of Green Hab operations: GreenHab door screws tightened.

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report.

Copyright © The Mars Society. All rights reserved. | Main Site